The vertex form of the equation of a parabola is y=(x+5)^2+49. What is the standard form of the equation?
Found 2 solutions by stanbon, swincher4391:
Answer by stanbon(75887)

The vertex form of the equation of a parabola is y=(x+5)^2+49. What is the standard form of the equation?
-----
Standard Form: y = ax^2 + bx + c
---
Your Problem:
y = x^2 + 10x+25 + 49
-----
y = x^2 + 10x + 74
=======================
Cheers,
Stan H.

Answer by swincher4391(1107)

y = (x+5)^2 + 49
y = x^2+10x + 25 + 49
y = x^2 +10x + 74